OpenStudy (anonymous):

explain cosx=1/sqrt2 = pi/4

OpenStudy (anonymous):

see when i checked the unit circle i could not find an x that was equal to 1/square root of 2

OpenStudy (anonymous):

the answer in the says pi/4, 7pi/4 or 3pi/4 and 5pi/4

OpenStudy (anonymous):

i figure since cosx=x/r, i should be able to get the radians from finding an x on the unit circle that equals to 1/sqrt of 2

OpenStudy (anonymous):

no way does \(\frac{1}{\sqrt2}=\frac{\pi}{4}\) what you may mean is IF \(\cos(x)=\frac{1}{\sqrt2}\) THEN \(x=\frac{\pi}{4}\)

OpenStudy (anonymous):

yes sorry

OpenStudy (anonymous):

oooh i see the problem

OpenStudy (anonymous):

\[\sqrt{\frac{1}{2}}=\frac{1}{\sqrt2}=\frac{\sqrt2}{2}\]

OpenStudy (anonymous):

?? how did the last step work

OpenStudy (anonymous):

you probably saw the third one on the unit circle and didn't recognize it as the second one

OpenStudy (anonymous):

rationalize the denominator

OpenStudy (anonymous):

\[\frac{1}{\sqrt2}\times \frac{\sqrt2}{\sqrt2}=\frac{\sqrt2}{2}\]

OpenStudy (anonymous):

oh i see how do you recognize those spots? or do you just need to have a feel for it?

OpenStudy (anonymous):

they are all the square root of one half

OpenStudy (anonymous):

this is the only one you will see in both forms probably

OpenStudy (anonymous):

you will get used to it

OpenStudy (anonymous):

ok, thank you!

OpenStudy (anonymous):

yw

OpenStudy (anonymous):

soryr first time using this, how do i say the question has been asnwered?

OpenStudy (anonymous):

sorry it took a while, site crashed on me

OpenStudy (anonymous):

you can "close' the question up top